# Quantum conditional mutual information as a probe of measurement-induced entanglement phase transitions

### v1

@@ I Introduction
Measurement-induced phase transitions (MIPTs) in quantum circuits composed of unitary dynamics and local projective measurements have attracted significant interest in recent years [1, 2, 3, 4, 5]. These transitions separate two distinct entanglement phases: a volume-law phase, in which the entanglement entropy of a subsystem scales proportionally with its volume, and an area-law phase, where it scales with the boundary area of the subsystem. Such transitions emerge from the competition between unitary dynamics, which generally generate entanglement among qubits, and local projective measurements, which tend to disentangle the measured qubits from the rest of the system. Remarkably, measurement-induced entanglement phase transitions [6, 7], as well as other nonequilibrium phase transitions [8], have been demonstrated experimentally on real quantum devices. These experimental breakthroughs have been made possible either by employing mid-circuit measurement and reset functionalities [7, 8], or by using a reference qubit initially maximally entangled with the monitored system as a probe for the entanglement phase transition [6, 9].
To numerically detect the measurement-induced entanglement phase transitions, the entanglement entropy itself is not always a suitable quantity, particularly due to its emergent logarithmic dependence on subsystem size at criticality in (1+1)$d$ systems [4, 2]. Indeed, performing finite-size scaling with entanglement entropy requires subtracting its value at an estimated critical point (e.g., a critical measurement rate) to eliminate the logarithmic dependence [4], which introduces an additional source of uncertainty. As an alternative, the tripartite mutual information (TMI) [10], defined as $I_{3}(A,B,C)=S_{A}+S_{B}+S_{C}-S_{A\cup B}-S_{B\cup C}-S_{A\cup C}-S_{A\cup B\cup C}$, where $S_{X}$ is the entanglement entropy of subsystem $X$, has been used to prove MIPTs [9, 11]. With an appropriately chosen tripartiton $A,B,$ and $C$, this linear combination of entanglement entropies effectively distinguishes volume-law and area-law phases without being affected by the logarithmic dependence at criticality. This approach has proven successful in systems with variable-range (power-law decaying) interactions [12, 13, 14, 15] as well as in systems with sparse nonlocal interactions [16]. While the effectiveness of the TMI is empirically established, it remains unclear–at least to our knowledge–why the logarithmic contribution is suppressed, especially given that $I_{3}(A,B,C)$ involves an odd number (seven) of entanglement entropy terms.
In this paper, we propose that another linear combination of entanglement entropies, of the form [17],
@@
with an appropriately chosen partition $A,B,$ and $C$ of the system, serves as an effective and pertinent probe of MIPTs. The quantity $I(A:C|B)$ is known as the quantum conditional mutual information (QCMI), and has been recently employed to quantify entanglement growth in stabilizer circuits composed of nearest-neighbor two-qubit Clifford gates [18]. Here, we demonstrate that the QCMI is also well-suited for probing measurement-induced entanglement phase transitions, through numerical simulations of one-dimensional variable-range ($1/r^{\alpha}$) Clifford circuits interspersed with local projective measurements. While similar models have been studied previously [13, 15], our circuit definition differs slightly in detail. Through careful finite-size-scaling analyses of $I(A:C|B)$, we find that no phase transition occurs for $\alpha\lesssim 2$, whereas for $2\lesssim\alpha\lesssim 3$, measurement-induced entanglement phase transitions emerge with critical exponents distinct from those in nearest-neighbor circuits and in the regime $3\lesssim\alpha$. These findings are qualitatively consistent with earlier studies on monitored circuits with variable-range interactions, not restricted to Clifford circuits [12, 13, 14, 15]. Furthermore, using crossing-point analysis of $I(A,C|B)$, we extract the coefficient $\tilde{c}$ of the emergent logarithmic scaling of the entanglement entropy at criticality in the thermodynamic limit. For the shortest-range interacting case studied, we find $\tilde{c}=1.519(3)$, which is significantly smaller than previous estimates obtained in monitored nearest-neighbor random Clifford circuits [2, 11, 19].
@@ II.1 Stabilizer circuit
We consider a one-dimensional chain of $L$ qubits with periodic boundary conditions, assuming that $L$ is even. Let $|\phi_{\zeta}(t)\rangle$ denote the pure quantum state of the entire system of $L$ qubits at discrete time step (simply referred to as “time”) $t$, initialized as $|\phi_{\zeta}(t=0)\rangle=|0\rangle^{\otimes L}$. Here, $\zeta$ indexes different circuit realizations characterized by random seeds. Time $t$ is incremented by one after each application of $L$ two-qubit Clifford gates, followed by projective measurements on each qubit along the Pauli-$Z$ axis with probability $p$ (see dashed horizontal lines in Fig. 1). Each two-qubit Clifford gate is randomly selected from the complete set of 11520 distinct two-qubit Clifford gates [20, 21], and applied to a randomly chosen pair of qubits $i$ and $j(\not{=}i)$ with probability
@@
where $r_{ij}$ is the distance between qubits $i$ and $j$ under periodic boundary conditions, defined as $r_{ij}=|i-j|$ for $|i-j|\leqslant L/2$ and $r_{ij}=L-|i-j|$ for $|i-j|>L/2$, ${\cal N}=\sum_{x=1}^{L/2}\frac{1}{x^{\alpha}}$ is the normalization constant, and $\alpha$ is a parameter that controls the interaction range of the two-qubit gates. We simulate circuit dynamics using the tableau algorithm [22] (see also Supplemental Material [23]), for system sizes ranging from $L=16$ to $2048$, and evolution times up to $t_{\rm max}=4096$. The parameters $\alpha$ and $p$ are varied in the ranges $1\leqslant\alpha\leqslant 4$ and $10\leqslant p<$ 1, respectively.
@@ II.2 Entanglement entropy and quantum conditional mutual information
where $\hat{\rho}_{X}(t)={\rm Tr}_{\overline{X}}\left[|\phi_{\zeta}(t)\rangle\langle\phi_{\zeta}(t)|\right]$ is the reduced density matrix of subsystem $X$, and $\bar{X}$ denotes its complement. To compute the entanglement entropy in stabilizer circuits, we follow the method described in Ref. [24] (see also Supplemental Material [23]). Note that when $p=1$, the entanglement entropy vanishes for all subsystems since the state $|\phi_{\zeta}(t)\rangle$ becomes a product state immediately after every measurement.
@@ II.3 Entanglement entropy at criticality
Motivated by the emergent conformal criticality observed in monitored Clifford and non-Clifford circuits with short-range interactions or hoppings [4, 2, 25, 26, 12, 13], we assume that the steady-state entanglement entropy $S_{X}=\frac{1}{N_{t}}\sum_{k=0}^{N_{t}-1}\langle\langle S_{X,\zeta}(t_{\rm min}+k\Delta_{S_{X},t})\rangle\rangle$ at criticality behaves as [27]
@@
where $l_{X}\equiv\frac{L}{\pi}\sin\frac{\pi|X|}{L}$ is the chord length of subsystem $X$ with $|X|$ denoting the number of qubits in $X$, $p_{\rm c}$ the critical value of the measurement rate, $c$ a universal coefficient of the logarithmic term (often referred to as an effective central charge) [28], and $c^{\prime}$ a nonuniversal constant. The autocorrelation time $\Delta_{S_{X},t}$ for the steady-state entanglement entropy is estimated in the same manner as $\Delta_{I,t}$, and exhibits similar behavior. Notice that the coefficient $c$ here is not equivalent to the central charge in conformal field theory. For example, in the case of the percolation universality class, the central charge is known to be zero [28, 29]. Note also that in the deep volume-law phase ($p\ll p_{c}$), the entanglement entropy scales extensively as $S_{X}\sim|X|$, while in the area-law phase ($p\gg p_{\rm c}$), it scales with the boundary, $S_{X}\sim|\partial X|$, where $|\partial X|$ denotes the number of qubits on the boundary of subsystem $X$.
Figure 2: Two distinct partitionings of the $L$-qubit system into subsystems $A$ (green), $B$ (white), $C$ (red), and $\overline{A\cup B\cup C}$ (gray). Each circle represents a qubit, and the number of qubits in each segment is indicated. Partition (a) is used for the QCMI in Eq. (7), while partition (b) is employed for the QCMI in Eq. (8).
@@ II.4 Quantum conditional mutual information at criticality
indicating that the nonuniversal constant $c^{\prime}$ cancels out in $I(A:C|B)$ at criticality. (ii) To eliminate the dependence of $I(A:C|B)$ on the total system size $L$ at criticality, we choose the subsystem sizes such that $|A\cup B|$, $|B\cup C|$, $|B|$, and $|A\cup B\cup C|$ are all proportional to $L$. (iii) To capture the volume-law entanglement phase, the subsystem sizes are chosen such that $L_{A\cup B}+L_{B\cup C}\neq L_{B}+L_{A\cup B\cup C}$, because for $p\ll p_{\rm c}$, the QCMI behaves as $I(A:C|B)\sim L_{A\cup B}+L_{B\cup C}-L_{B}-L_{A\cup B\cup C}$ , where $L_{X}\equiv\min(|X|,L-|X|)$ is the smaller of the number of qubits in $X$ and its complement $\overline{X}$. We note that in the area-law phase $p\gg p_{\rm c}$, the QCMI behaves as $I(A:C|B)\sim 0$, regardless of the partitioning. Moreover, the inequality $I(A:C|B)\geqslant 0$ always holds due to the strong subadditivity of von Neumann entanglement entropy [30]. Thus, as long as the partitioning satisfies conditions (i)-(iii), the QCMI $I(A:C|B)$ serves as a meaningful probe of the measurement-induced entanglement phase transition. Importantly, condition (iii) is independent of the assumption of conformal criticality in Eq. (5), implying that $I(A:C|B)$ can detect the entanglement phase transition even in the absence of emergent conformal invariance.
Specifically, we consider two types of partitioning for the same system, in a spirit of Ref. [31]. The first partitioning sets $|A|=|B|=|C|=L/4$ [see Fig. 2(a)], for which we find
@@
Here, the subscripts $a$ and $b$ in $I_{a}$ and $I_{b}$ correspond to the partitionings illustrated in Figs. 2(a) and 2(b), respectively. It should be emphasized that Eqs. (6)–(8) are valid only under the assumption of logarithmic scaling of the entanglement entropy at criticality, as expressed in Eq. (5). As we show in the next section, our numerical results strongly support the validity of this assumption for $\alpha\gtrsim 3$.
@@ III.1 Phase diagram
Before presenting a detailed analysis, we first show the global phase diagram of the stabilizer circuits with variable-range two-qubit Clifford gates as a function of $\alpha$ and $p$ in Fig. 3. Using the QCMI as a probe, we accurately determine the phase boundary separating the area-law and volume-law entanglement phases for $\alpha>2$. In what follows, we analyze the critical properties of the system along this phase boundary. Additionally, we confirm that no area-law phase exists for $\alpha\lesssim 2$, which is consistent with the previous analytical arguments in the high-measurement-rate regime [13, 15].
@@ III.2 Quantum conditional mutual information near criticality
Let us denote the QCMI as $I_{\gamma}(p,L)$, where $\gamma=a$ and $b$ correspond to the two distinct partitioning schemes shown in Figs. 2(a) and 2(b), respectively. At the critical point $p=p_{\mathrm{c}}$, the QCMI becomes independent of the system size [see Eqs. (7) and (8)], which suggests that the crossing-point analysis [32] can be effectively used to estimate both $p_{\mathrm{c}}$ and the correlation-length critical exponent $\nu$. Indeed, as shown in Fig. 4, the curves of $I_{\gamma}(p,L)$ for different $L$ appear to cross at a single point for sufficiently large system sizes. The estimated critical points $p_{\mathrm{c}}$ are nearly identical for both partitionings $\gamma=a$ and $b$, for each value of $\alpha$. This already demonstrates, at least in part, that the QCMI serves as a reliable probe to detect the critical point. For smaller values of $\alpha$, deviations from the size-independent behavior are more pronounced in smaller systems [see Figs. 4(b) and 4(d)]. Nevertheless, well-established methods for handling corrections to scaling have been developed and documented in the literature [33]. Interestingly, we find that corrections to scaling becomes more subtle for larger $\alpha$. As discussed in Sec. III.4, nonmonotonic finite-size behavior of the crossing points [see Figs. 8(a) and 9(a)] emerges in the vicinity of the critical point, suggesting that multiple correction terms may need to be considered. A similar issue was only recently resolved in the study of the Néel-paramagnetic quantum phase transition in dimerized Heisenberg models, where the known value of the correlation-length exponent $\nu$ from the expected O(3) universality class was crucial for the analysis [34]. In our case, however, neither the critical exponent $\nu$ nor the universality class is known a priori, necessitating the use of an alternative approach, as described in the following.
Figure 4: QCMI $I_{\gamma}(p,L)$ for different partitionings and interaction ranges. Panels (a) and (b) show results for $\gamma=a$, while (c) and (d) correspond to $\gamma=b$. The interaction-range parameter is $\alpha=3.5$ for (a) and (c), and $\alpha=2.5$ for (b) and (d). Vertical dashed lines indicate the critical points estimated from the data-collapse analyses in Figs 6 and 7.

### v2

@@ I Introduction
Measurement-induced phase transitions (MIPTs) in quantum circuits composed of unitary dynamics and local projective measurements have attracted significant interest in recent years [Li-PRB2018, Li-PRB2019a, Chan-PRB2019, Skinner-PRX2019, Fisher-ARCMP2023]. These transitions separate two distinct entanglement phases: a volume-law phase, in which the entanglement entropy of a subsystem scales proportionally with its volume, and an area-law phase, where it scales with the boundary area of the subsystem. Such transitions emerge from the competition between unitary dynamics, which generally generate entanglement among qubits, and local projective measurements, which tend to disentangle the measured qubits from the rest of the system. Remarkably, measurement-induced entanglement phase transitions [Noel-NatPhys2022, Koh-NP2023], as well as other nonequilibrium phase transitions [Chertkov_NatPhys2023], have been demonstrated experimentally on real quantum devices. These experimental breakthroughs have been made possible either by employing mid-circuit measurement and reset functionalities [Koh-NP2023, Chertkov_NatPhys2023], or by using a reference qubit initially maximally entangled with the monitored system as a probe for the entanglement phase transition [Noel-NatPhys2022, Gullans-PRX2020].
To numerically detect the measurement-induced entanglement phase transitions, the entanglement entropy itself is not always a suitable quantity, particularly due to its emergent logarithmic dependence on subsystem size at criticality in (1+1)$d$ systems [Skinner-PRX2019, Li-PRB2019a]. Indeed, performing finite-size scaling with entanglement entropy requires subtracting its value at an estimated critical point (e.g., a critical measurement rate) to eliminate the logarithmic dependence [Skinner-PRX2019], which introduces an additional source of uncertainty. As an alternative, the tripartite mutual information (TMI) [Kitaev-PRL2006], defined as $I_{3}(A,B,C)=S_{A}+S_{B}+S_{C}-S_{A\cup B}-S_{B\cup C}-S_{A\cup C}-S_{A\cup B\cup C}$, where $S_{X}$ is the entanglement entropy of subsystem $X$, has been used to prove MIPTs [Gullans-PRX2020, Zabalo-PRB2020]. With an appropriately chosen tripartiton $A,B,$ and $C$, this linear combination of entanglement entropies effectively distinguishes volume-law and area-law phases without being affected by the logarithmic dependence at criticality. This approach has proven successful in systems with variable-range (power-law decaying) interactions [Minato-PRL2022, Block-PRL2022, Muller-PRL2022, Sahu-PRB2022] as well as in systems with sparse nonlocal interactions [Hashizume-PRR2022]. While the effectiveness of the TMI is empirically established, it remains unclear–at least to our knowledge–why the logarithmic contribution is suppressed, especially given that $I_{3}(A,B,C)$ involves an odd number (seven) of entanglement entropy terms.
In this paper, we propose that another linear combination of entanglement entropies, of the form [Levin-PRL2006],
@@
with an appropriately chosen partition $A,B,$ and $C$ of the system, serves as an effective and pertinent probe of MIPTs. The quantity $I(A:C|B)$ is known as the quantum conditional mutual information (QCMI) [Cerf-PRL1997, Cerf-PhysicaD1998], and has been recently employed to quantify measurement-induced topological phase transitions [Lavasani-NatPhys2021, Lavasani-PRL2021, Kells-SciRep2023], and entanglement growth in stabilizer circuits composed of nearest-neighbor two-qubit Clifford gates [Sang-PRXQ2023]. The QCMI has also been used to characterize one-dimensional topological insulators and superconductors at equilibrium [Maiellaro-PRR2022, Maiellaro-PRB2023].
Here, we demonstrate that the QCMI is also well-suited for probing measurement-induced entanglement phase transitions, through numerical simulations of one-dimensional variable-range ($1/r^{\alpha}$) Clifford circuits interspersed with local projective measurements. While similar models have been studied previously [Block-PRL2022, Sahu-PRB2022], our circuit definition differs slightly in detail. We show that the QCMI is free from multiplicative and additive renormalization; at criticality, it becomes system-size independent under suitable partitioning of the system. This property makes the QCMI particularly useful for identifying critical points and extracting universal quantities much like the Binder ratio in conventional statistical physics. Notably, the value of the QCMI at emergent conformal criticality directly yields the universal coefficient of the logarithmic term. Through careful finite-size-scaling analyses of $I(A:C|B)$, we find that no phase transition occurs for $\alpha\lesssim 2$, whereas for $2\lesssim\alpha\lesssim 3$, measurement-induced entanglement phase transitions emerge with critical exponents distinct from those in nearest-neighbor circuits and in the regime $3\lesssim\alpha$. These findings are qualitatively consistent with earlier studies on monitored circuits with variable-range interactions, not restricted to Clifford circuits [Minato-PRL2022, Block-PRL2022, Muller-PRL2022, Sahu-PRB2022]. Furthermore, using crossing-point analysis of $I(A,C|B)$, we extract the coefficient $\tilde{c}$ of the emergent logarithmic scaling of the entanglement entropy at criticality in the thermodynamic limit. For the shortest-range interacting case studied, we find $\tilde{c}=1.519(3)$, which is significantly smaller than previous estimates obtained in monitored nearest-neighbor random Clifford circuits [Li-PRB2019a, Zabalo-PRB2020, Sierant-PRB2022].
@@ II.1 Stabilizer circuit
We consider a one-dimensional chain of $L$ qubits with periodic boundary conditions, assuming that $L$ is even. Let $|\phi_{\zeta}(t)\rangle$ denote the pure quantum state of the entire system of $L$ qubits at discrete time step (simply referred to as “time”) $t$, initialized as $|\phi_{\zeta}(t=0)\rangle=|0\rangle^{\otimes L}$. Here, $\zeta$ indexes different circuit realizations characterized by random seeds. Time $t$ is incremented by one after each application of $L$ two-qubit Clifford gates, followed by projective measurements on each qubit along the Pauli-$Z$ axis with probability $p$ (see dashed horizontal lines in Fig. 1). Each two-qubit Clifford gate is randomly selected from the complete set of 11520 distinct two-qubit Clifford gates [Corcoles-PRA2013, Richter-PRR2023], and applied to a randomly chosen pair of qubits $i$ and $j(\not{=}i)$ with probability
@@
where $r_{ij}$ is the distance between qubits $i$ and $j$ under periodic boundary conditions, defined as $r_{ij}=|i-j|$ for $|i-j|\leqslant L/2$ and $r_{ij}=L-|i-j|$ for |ij|>L/2|i-j|>L/2, ${\cal N}=\sum_{x=1}^{L/2}\frac{1}{x^{\alpha}}$ is the normalization constant, and $\alpha$ is a parameter that controls the interaction range of the two-qubit gates. We simulate circuit dynamics using the tableau algorithm [Aaronson-PRA2004] (see also Supplemental Material [suppl]), for system sizes ranging from $L=16$ to $2048$, and evolution times up to $t_{\rm max}=4096$. The parameters $\alpha$ and $p$ are varied in the ranges $1\leqslant\alpha\leqslant 4$ and $0\leqslant p<1$, respectively.
@@ II.2 Entanglement entropy and quantum conditional mutual information
where $\hat{\rho}_{X}(t)={\rm Tr}_{\overline{X}}\left[|\phi_{\zeta}(t)\rangle\langle\phi_{\zeta}(t)|\right]$ is the reduced density matrix of subsystem $X$, and $\bar{X}$ denotes its complement. To compute the entanglement entropy in stabilizer circuits, we follow the method described in Ref. [Nahum-PRX2017] (see also Supplemental Material [suppl]). Note that when $p=1$, the entanglement entropy vanishes for all subsystems since the state $|\phi_{\zeta}(t)\rangle$ becomes a product state immediately after every measurement.
@@ II.3 Entanglement entropy at criticality
Motivated by the emergent conformal criticality observed in monitored Clifford and non-Clifford circuits with short-range interactions or hoppings [Skinner-PRX2019, Li-PRB2019a, Fuji-PRB2020, Alberton-PRL2021, Minato-PRL2022, Block-PRL2022], we assume that the steady-state entanglement entropy $S_{X}=\frac{1}{N_{t}}\sum_{k=0}^{N_{t}-1}\langle\langle S_{X,\zeta}(t_{\rm min}+k\Delta_{S_{X},t})\rangle\rangle$ at criticality behaves as [Calabrese-JPAMT2009]
@@
where $l_{X}\equiv\frac{L}{\pi}\sin\frac{\pi|X|}{L}$ is the chord length of subsystem $X$ with $|X|$ denoting the number of qubits in $X$, $p_{\rm c}$ the critical value of the measurement rate, $c$ a universal coefficient of the logarithmic term (often referred to as an effective central charge) [Jian-PRB2020], and $c^{\prime}$ a nonuniversal constant. The autocorrelation time $\Delta_{S_{X},t}$ for the steady-state entanglement entropy is estimated in the same manner as $\Delta_{I,t}$, and exhibits similar behavior. Notice that the coefficient $c$ here is not equivalent to the central charge in conformal field theory. For example, in the case of the percolation universality class, the central charge is known to be zero [Jian-PRB2020, Oshima-PRB2023]. Note also that in the deep volume-law phase ($p\ll p_{c}$), the entanglement entropy scales extensively as $S_{X}\sim|X|$, while in the area-law phase ($p\gg p_{\rm c}$), it scales with the boundary, $S_{X}\sim|\partial X|$, where $|\partial X|$ denotes the number of qubits on the boundary of subsystem $X$.
Figure 2: Two distinct partitions of the $L$-qubit system into subsystems AA (green), $B$ (white), $C$ (red), and $\overline{A\cup B\cup C}$ (gray). Each circle represents a qubit, and the number of qubits in each segment is indicated. Partition (a) is used for the QCMI in Eq. (7), while partition (b) is employed for the QCMI in Eq. (8).
@@ II.4 Quantum conditional mutual information at criticality
indicating that the nonuniversal constant $c^{\prime}$ cancels out in $I(A:C|B)$ at criticality. (ii) To eliminate the dependence of $I(A:C|B)$ on the total system size $L$ at criticality, we choose the subsystem sizes such that $|A\cup B|$, $|B\cup C|$, $|B|$, and $|A\cup B\cup C|$ are all proportional to $L$. (iii) To capture the volume-law entanglement phase, the subsystem sizes are chosen such that $L_{A\cup B}+L_{B\cup C}\neq L_{B}+L_{A\cup B\cup C}$, because for $p\ll p_{\rm c}$, the QCMI behaves as $I(A:C|B)\sim L_{A\cup B}+L_{B\cup C}-L_{B}-L_{A\cup B\cup C}$ , where $L_{X}\equiv\min(|X|,L-|X|)$ is the smaller of the number of qubits in $X$ and its complement $\overline{X}$. We note that in the area-law phase $p\gg p_{\rm c}$, the QCMI behaves as $I(A:C|B)\sim 0$, regardless of the partitioning. Moreover, the inequality $I(A:C|B)\geqslant 0$ always holds due to the strong subadditivity of von Neumann entanglement entropy [Lieb-PRL1973]. Thus, as long as the partitioning satisfies conditions (i)-(iii), the QCMI $I(A:C|B)$ serves as a meaningful probe of the measurement-induced entanglement phase transition. Importantly, condition (iii) is independent of the assumption of conformal criticality in Eq. (5), implying that $I(A:C|B)$ can detect the entanglement phase transition even in the absence of emergent conformal invariance.
Specifically, we consider two types of partitioning for the same system, in a spirit of Ref. [Furukawa-PRL2009]. The first partitioning sets $|A|=|B|=|C|=L/4$ [see Fig. 2(a)], for which we find
@@
Here, the subscripts $a$ and $b$ in $I_{a}$ and $I_{b}$ correspond to the partitions illustrated in Figs. 2(a) and 2(b), respectively. It should be emphasized that Eqs. (6)–(8) are valid only under the assumption of logarithmic scaling of the entanglement entropy at criticality, as expressed in Eq. (5). As we show in the next section, our numerical results strongly support the validity of this assumption for $\alpha\gtrsim 3$.
@@ III.1 Phase diagram
Before presenting a detailed analysis, we first show the global phase diagram of the stabilizer circuits with variable-range two-qubit Clifford gates as a function of $\alpha$ and $p$ in Fig. 3. Using the QCMI as a probe, we accurately determine the phase boundary separating the area-law and volume-law entanglement phases for $\alpha>2$. In what follows, we analyze the critical properties of the system along this phase boundary. Additionally, we confirm that no area-law phase exists for $\alpha\lesssim 2$, which is consistent with the previous analytical arguments in the high-measurement-rate regime [Block-PRL2022, Sahu-PRB2022].
@@ III.2 Quantum conditional mutual information near criticality
Let us denote the QCMI as $I_{\gamma}(p,L)$, where $\gamma=a$ and $b$ correspond to the two distinct partitioning schemes shown in Figs. 2(a) and 2(b), respectively. At the critical point $p=p_{\mathrm{c}}$, the QCMI becomes independent of the system size [see Eqs. (7) and (8)], which suggests that the crossing-point analysis [Luck-PRB1985] can be effectively used to estimate both $p_{\mathrm{c}}$ and the correlation-length critical exponent $\nu$. Indeed, as shown in Fig. 4, the curves of $I_{\gamma}(p,L)$ for different $L$ appear to cross at a single point for sufficiently large system sizes. The estimated critical points $p_{\mathrm{c}}$ are nearly identical for both partitions $\gamma=a$ and $b$, for each value of $\alpha$. This already demonstrates, at least in part, that the QCMI serves as a reliable probe to detect the critical point. For smaller values of $\alpha$, deviations from the size-independent behavior are more pronounced in smaller systems [see Figs. 4(b) and 4(d)]. Nevertheless, well-established methods for handling corrections to scaling have been developed and documented in the literature [Shao-S2016]. Interestingly, we find that corrections to scaling becomes more subtle for larger $\alpha$. As discussed in Sec. III.4, nonmonotonic finite-size behavior of the crossing points [see Figs. 8(a) and 9(a)] emerges in the vicinity of the critical point, suggesting that multiple correction terms may need to be considered. A similar issue was only recently resolved in the study of the Néel-paramagnetic quantum phase transition in dimerized Heisenberg models, where the known value of the correlation-length exponent $\nu$ from the expected O(3) universality class was crucial for the analysis [Ma-PRL2018]. In our case, however, neither the critical exponent $\nu$ nor the universality class is known a priori, necessitating the use of an alternative approach, as described in the following.
Figure 4: QCMI $I_{\gamma}(p,L)$ for different partitions and interaction ranges. Panels (a) and (b) show results for $\gamma=a$, while (c) and (d) correspond to $\gamma=b$. The interaction-range parameter is $\alpha=3.5$ for (a) and (c), and $\alpha=2.5$ for (b) and (d). Vertical dashed lines indicate the critical points estimated from the data-collapse analyses in Figs 6 and 7.